ERRATA TO The Joy of TEX PRIOR TO AMS-TEX 2.0

This list of corrections to The Joy of TEX, 1986 edition, includes all known corrections that preceded the release of AMS-TEX Version 2.0. Reprints with corrections may already incorporate some or all of these changes.

The printing date of each copy of The Joy of TEX is identified on the reverse of the title page. The list below will permit you to determine which corrections have not already been incorporated in your copy of The Joy of TEX.

First printing, 1986  all changes
Second printing with corrections, 1986  changes after 11/25/86
Third printing with corrections, 1987  changes after 5/12/87

For differences between earlier versions of AMS-TEX and Version 2.0, see the User’s Guide to AMS-TEX 2.0. The second edition of The Joy of TEX, 1990, contains all changes in this list as well as new material for AMS-TEX 2.0.

(This errata list was last updated 15 October 89.)

Page 12, line 12 (11/11/86)

What output is produced by \$\␣\␣1.00 and by \$␣\␣1.00?

Page 22, line 28 (11/24/86)

will be some surprises in it—so you should go pick it up as soon as possible.

Page 26, line 9 (10/15/89)

uptight when you encounter an error message, because TEX can always be coaxed

Page 39, line 4 (10/15/89)

words as evenly as possible. But everyone knows that such bland perfection isn’t

Page 39, line −4 (12/12/89)

allowed here also, to accommodate threesomes, foursomes, and even more perverse

Page 44, line −10 (12/12/89)

their own papers might prefer to leave these details to someone else, and even

Page 81, line 13 (10/25/89)

But don’t use \, before an expression like \frac{dy}{dx} or before the dx in dy/dx.

Page 88, line −5 (5/11/87)

We derive the quadratic formula by “completing the square”:

Typeset by AMS-TEX

Page 90, line −4 (10/15/89) to the old style that they may be discomforted by the “improvements”.

Page 99, lines 15–16 (8/6/86)

       $\varinjlim$                           \underrightarrow{lim}
       $\varprojlim$                          \underleftarrow{lim}

Page 108, line 11 (11/11/86)

         &=(a+b)(a+b)^n=(a+b)

Page 109, line 6 (12/12/89) when tags are set on the right. What input do you think you should use?

Page 109, line −14 (10/15/89) so that the =\bigl[ is aligned with the invisible \qquad. Notice, again, that such

Page 113, line 1 (4/10/86)

And there’s \bmatrix...\endbmatrix to get brackets \left[...\right] around

Page 127, line 11 (7/13/87) If you’re an experienced mathematical typist you’ve probably already begun to

Page 129, lines 14–15 (10/15/89) with things like (x_{1}, …, x_{m}), (y_{1}, …, y_{n+1}) as well. Explain how to define

\vector so that we can type these as $\vector xm$ and $\vector y{n+1}$.

Page 129, last 3 lines (10/15/89) In Exercise 19.20 we defined \vector so that $\vector xn$ produces (x_{1}, …, x_{n}), etc. But perhaps you don’t like this, perhaps you’d prefer to type $\vector nx$, with the ‘n’ first, and the ‘x’ second. How can you arrange this?

Page 131, lines 10–11 (10/15/89) How would you \define the control sequence \vector so that you type $\vector x,n.$ to get (x_{1}, …, x_{n}), and $\vector y,m+1.$ to get (y_{1}, …, y_{m+1}), etc.

Page 144, line 16 (10/15/89) This command is “global”—it affects everything that follows, even if it is in-

Page 162, line −6 (5/11/87) if you typed \footnote""{...} then you would get no marker at all, just a note

Page 171, line −7 (10/15/89) too much, and only \linebreak will force TEX to overcome its reluctance.

Page 176, line 4 (12/12/89) about it, and an & is tolerated only in special situations. So you should remember

Page 179, line 4 (10/15/89) change its position on the 8\frac{1}{2} by 11 sheet of paper. Typing

Page 180, lines 5–6 (10/15/89)

          &=f'(x) = \frac1{2\sqrt x}\qquad
       \foldedtext\foldedwidth{2in}{for some $x$ in $(k, k+1)$,

Page 181, line −4 (10/15/89) should be included at the end of that displayed formula.

Page 182, line −2 (12/12/89) argument” feature of \roster (again compare with footnote). If you type

Page 186, line 13 (10/15/89) commands are “global”—they affect everything that follows even if used in a group

Page 189, line 21 (12/12/89) will first be divided into lines of a certain length (3 inches less than the width

Page 195, lines 4, 11 (7/13/87)

Change  “In addition to”  to  “First we have”.

Page 195, line −1 (12/12/89)

       ... in a bibliography''.

Page 202, line −6 (12/12/89) If ‘etc.’ were typed instead of ‘etc\.’ there would be a larger space after the

Page 208, line 12 (12/12/89) it does in ordinary text.

Page 210, line 4 (12/12/89) you’ll get the two equations a + b = c and A + B = C displayed separately.

Page 212, line 6 (12/12/89) If you press ⟨carriage-return⟩, TEX will continue merrily, and you will get a^{b}^{c}

Page 218, line −6 (7/13/87) Of course, you weren’t supposed to anticipate such after-the-fact corrections.

Page 222, answer to 14.11, line 1 (10/15/89)

       We derive the quadratic formula by

Page 229, answer to 15.19, lines 2–3 (10/15/89)

       $\operatorname{\text{\sl SO}}(n)$  SO(n)
       $\operatorname{\text{\bf SO}}(n)$  SO(n)

Page 230, answer to 16.3, lines 6–9 (10/25/89)

to suppress any extra space that TEX might put in. (Actually, ...\tag{$**$}$$ happens to work correctly, but ...\tag{$***$}$$ would give the tag (∗ ∗ ∗); rather than worrying about why this happens, just type ...\tag{${*}{*}$}$$

and ...\tag{${*}{*}{*}$}$$ to be on the safe side.)

Page 230, answer to 16.4, line 3 (7/13/87)

       Q^l&=Q_1\biggl\{\sum_k(-1)^k(PQ_1-I)^k\biggr\}

Page 230, answer to 16.4, line 6 (10/25/89)

           Q_1\tag 1{${}_r$}

Page 231, answer to 16.6 (10/25/89)

Line 2:

       \align \alpha_4&=\sqrt{\dfrac12}\\

Line 6:

       \text{etc.}

Page 233, answer to 17.4, line 6 (5/13/86)

               \dots, $b_{3k}$.}\endmultline

Page 234, answer to 18.4 (5/13/86)

Line 6:

       \pmatrix \format\r&\quad\r\\

Line 10:

       =\pmatrix \format\r&\quad\r\\

Page 239, answer to 19.13 (10/15/89)

       \define\vector#1{(#1_1,\dots,#1_n)}

and then use $\vector x to get (x_{1}, …, x_{n}) and $\vector y$ to get (y_{1}, …, y_{n}), etc.

Page 240, answer to 19.14 (10/15/89)

       $\vector\alpha$ and $\vector{x'}$.

Page 240, answer to 19.15 (10/15/89) 19.15. You can get (x′_{1}, …, x′_{n}) by typing $\vector{{x'}}$; now the argument is {x'} and {x'}_1 gives x′_{1}, etc. On the other hand, you can’t get the formula (x_{1}′, …, x_{n}′) using \vector—you’d just have to type it out in full.

Page 240, answer to 19.20 (10/15/89)

       \define\vector#1#2{(#1_1,\dots,#1_{#2})}

Page 242, answer to 19.23 (10/15/89)

       \define\vector#1#2{(#2_1,\dots,#2_{#1})}

Although #1 and #2 must appear in that order after the \define\vector, they can appear in any order within the definition itself.

Page 242, answer to 19.24 (10/15/89)

       \define\vector#1,#2.{(#1_1,\dots,#1_{#2})}

Page 242, answer to 19.27, line 1 (7/13/87) 19.27. This is a perfectly acceptable \define, but you are not defining a new

Page 251, line 1 (10/25/89) is supplied as a synonym for \thickspace. In plain, the thick space \; can

Page 252, line 6 (11/11/86)

       $f''^2$

Page 261, after line 12 (6/22/87)

  Add  ≂ \eqsim

Page 261, line 15 (6/22/87)

  Change  ≇ \napprox  to  ≇ \ncong

Page 262, line 15 (11/14/86)

  Change  ð \thorn  to  ð \eth

Page 264, line 1 (11/11/86) Appendix G: {TEX Users}

Page 265, line 6 (11/11/86) you might want to look back at Appendix G. Perhaps someone in TUG has

Page 275, column 1 (11/14/86)

Add entry \eth (ð), 262 Page 279, column 1 (12/12/89)

\lesssim (≲), 260

Page 281, column 1 (6/22/86) Remove entry for \napprox Add entry

\ncong (≇), 261

Page 284, column 1 (12/12/89) \Psi (Ψ), 255

Page 288, column 2 (11/14/86) Delete entry for \thorn